Principle of mathematical induction to prove 1+2+2²+2³+....+2^n-1=2^n-1


1
Expert's answer
2022-06-06T11:27:56-0400

Let "P(n)" be the proposition that for the first "n" positive integers


"1+2+2^2+2^3+...+2^{n-1}=2^n-1"

Basis Step:

"P(1)"is true, because "1=2^1-1."

Inductive Step:

For the inductive hypothesis we assume that "P(k)" holds for an arbitrary

positive integer "k." That is, we assume that


"1+2+2^2+2^3+...+2^{k-1}=2^k-1"

Under this assumption, it must be shown that "P(k + 1)" is true, namely, that


"1+2+2^2+2^3+...+2^{k-1}+2^{k+1-1}=2^{k+1}-1"

is also true.


"1+2+2^2+2^3+...+2^{k-1}+2^{k+1-1}"

"=2^k-1+2^k=2(2^k)-1"

"=2^{k+1}-1"

This last equation shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all positive integers "n." That is, we have proved that

"1+2+2^2+2^3+...+2^{n-1}=2^n-1"

for all positive integers "n."